How many committees of four can be selected from a group of seven if the same person must be included in each committee?
Answer by Edwin McCravy(20055)

How many committees of four can be selected from a group of seven if the same person must be included in each committee?
Suppose the person who must be in each committee is named Bob.

The problem amounts to selecting 3 from the 6 people besides Bob
to put in a committee with Bob.  So the answer is "6 things choose 3".

6C3 = (6*5*4)/(3*2*1) = 120/6 = 30

Edwin